# How Good is Your Wikipedia?

### v1

@@ 5.2 Language adaptation
Table 6: XLM-R base, XLM-R with continued pretraining on unseen languages with raw_wiki, +primary and +heuristic partitions fine-tuned on SIB200. Results are averaged over 5 runs, and languages are grouped by tier.
Results are in Table 6, with the trends being slightly different from the trends seen with the monolingual models in Table 5. While all languages (except mni) benefit from language adaption, the effect of quality filtering is dependent upon the underlying data distribution. Tier 2 and Tier 3 benefit the most from primary filtering, with ace achieving an 8% increase. Tier 1 and Tier 4, on the other hand, do not see much performance improvement with filtering on average. Notably, mni suffers a staggering 16% performance drop after language adaption, raising many questions about the quality of that Wikipedia.
6 Discussion
6.1 Effect of Quality Filtering
We look for a correlation between data pruning and accuracy with a Spearman’s correlation test between the amount of data removed in the primary filtering step, and the accuracy difference between raw_wiki and +primary. There is no significant correlation, which means that data quality may not have a direct significant impact on performance. Findings from van Noord et al. (2024) and Artetxe et al. (2022) also echo this result. This means that there is no significant decrease in performance when removing much of the data. With results in Figure 6 we saw a reduction in data size by 50% with little impact on performance. Thus, we recommend data quality filtering not as a means for improving performance, but as a means for efficient training and achieving good performance with less compute.
Overall, the primary filtering step with script filtering and deduplication is beneficial for most Wikipedias. Further heuristic filtering, on the other hand, yields mixed results. This can be dependent upon the underlying distribution of the Wikipedia, since Wikipedias in Tier 2 and Tier 3 can benefit more from quality filtering than those in Tier 1, which already have good quality data distributions, and in Tier 4 with subpar data distributions (see Figure 4) that cannot be helped by filtering data any further.
6.2 Quality versus Size
Here we highlight the benefits of data quality over both data quantity and model size. While the discussions so far have mainly been centred around monolingual data quality, a comparison of results from Tables 5 and 6 highlight another salient aspect. When compared with the XLM-R model adapted to the same data (for e.g., monolingual raw_wiki versus XLM-R+raw_wiki), 23 monolingual models either perform on par or outperform XLM-R with a fraction of the data size and capacity. Our results for AfriSenti-Twitter are competitive with those reported by Muhammad et al. (2023) on larger multilingual models. This indicates that a smaller concentration of monolingual data trained with fewer parameters can yield better performance than a large multilingual corpus that includes the same data. Thus, low-resource languages benefit from smaller, more efficient models trained on a fraction of the compute. This is also seen with the models in Figure 6, which achieve competitive performance to models trained on the entire Wikipedia with essentially half the data.
6.3 Data Quality versus Domain
As discussed in §2.3, data quality is subjective and so domain may be a more useful consideration. We see a reflection of this when comparing the effect of data quality across tasks in Tables 2, 3 and 4. MasakhaNER favours the seemingly lower quality, unfiltered raw_wiki distribution. This may be due to the encyclopedic domain of Wikipedia lending itself very well to good performance on a task like NER. Quality filtering can remove the one-liner articles that describe named entities which can be beneficial for this task. Performance on classification tasks, on the other hand, benefits from the data quality filtering.
Thus, we push for a shift in perspective from seeking general definitions of data quality to domain and task specific investigations of what data quality means.
6.4 Data Quality versus Language
A push away from general definitions of data quality also leads us towards considering the effect of data quality on individual Wikipedias. The results discussed in this study on average remain consistent, but zooming in on individual languages reveals the large benefits of this process for some languages. We touched upon the results for specific languages in §5.1.1, showing that averages may not entirely reflect trends with individual Wikipedias. Therefore, we emphasize that there there is no one-size-fits-all data filtering approach. Data quality decisions for a corpus like Wikipedia containing multiple languages, even though they are all within the same domain, should be made on a per-language basis. This approach can be made more principled by leveraging typological features and databases to create more language-specific data filtering pipelines, which is a promising direction for future work.
7 Conclusion
With this study, we challenge the assumption of Wikipedia as a high quality resource in multilingual NLP, and provide a guide to using Wikipedia as a pretraining resource. We show that much of Wikipedia contains duplicate articles, foreign scripts, etc., which can be filtered out as a means to achieve good performance with less compute. For some languages, it can even lead to performance improvement. Overall, we advocate towards a shift in perspective from seeking general definitions of data quality to approaching quality issues from a domain or language specific lens.
8 Acknowledgements
Appendix A Pretraining Details
Hyperparameters for the pretraining and fine-tuning of monolingual models are in Table 7. We use the SentencePiece package Kudo and Richardson (2018) to train monolingual unigram tokenisers for each language. We train two tokenisers per language: one on the raw_wiki with no filtering, which is used to train the raw_wiki models; the second one on the data remaining after script filtering and deduplication, which we use to train all other models. We use an approximation with Zipf’s law Zipf (1949) on each Wikipedia to estimate the vocabulary size for each model, mentioned in Table 7. Hyperparameters for XLM-R language adaption are also in Table 7. For both cases, we checkpoint and use the best model on a left-out validation set to fine-tune after pretraining.
Tiny DeBERTa Hyperparameters Estimated Vocabulary Size
Pretraining Lang Vocab Lang Vocab
Model Type deberta ace 3505 mr 14708
Hidden Size 312 am 5234 nl 60011
Intermediate Size 1200 ary 4502 nso 2801
Attention Heads 12 awa 2190 ny 2472
Hidden Layers 4 azb 13996 om 3436
Hidden Act gelu ban 6624 pag 2170
Max Length 128 bm 1485 pcm 2595
Task MLM ceb 73124 rn 1446
Mask Probability 0.4 ckb 11451 ro 36329
Padding Strategy longest ee 1770 rw 5725
Grad. Accumulation Steps 4 fon 1426 shn 5784
Batch size 8 and 32 ha 13126 sn 5227
Learning Rate 1e-3 ht 10836 so 6117
Epochs 100 hu 46205 sw 12529
Finetuning id 41725 te 23175
Max Length 128 ig 11609 ti 1150
Epochs 20 kk 22496 tn 3783
Batch Size 8 ks 2248 ts 1817
Learning Rate 5e-5 lg 4528 tt 26047
Padding Strategy max_length lij 5462 tw 4745
XLMR Hyperparameters ln 2509 uk 62626
Language Adaption mai 4793 vi 43806
Model Type xlm-roberta mi 3451 war 25803
Task MLM min 14783 wo 3384
Max Length 512 mni 3351 yo 6128
Mask Probability 0.15 zu 4439
Epochs 3
Batch Size 8
Learning Rate 5.00e-5
Padding strategy Longest
Grad. Accumulation Steps 1
Finetuning
Max Length 512
Epochs 20
Batch Size 32
Learning Rate 5.00e-5
Padding Strategy max_length
Table 7: (1) Hyperparameters for monolingual pretraining and fine-tuning. While we used 100 epochs as a default, for Wikipedias that are too large (nl, ro, id, hu, uk, vi, ceb, war, tt, azb, min) we pretrain only for 2 epochs. Similarly, while we used a batch size of 8 for all models, for the larger Wikipedias we use a batch size of 32. (2) Hyperparameters for XLM-R language adaptation (based on Alabi et al. (2022)) and fine-tuning. (3) Estimated vocabulary size used to train monolingual tokenisers for each Wikipedia in our experimental pipeline.
Appendix B Results from partitioned models
Results for the partitioned models fine-tuned on SIB200 are in Table 8, MasakhaNER, MasakhaNEWS and AfriSenti-Twitter are in Table 9. Scores from this table were used to create the boxplot in Figure 6.
SIB200
Lang absolute relative entropy
low high low high low high
ha 0.77 0.77 0.75 0.71 0.72 0.77
ig 0.76 0.76 0.74 0.75 0.73 0.77
sn 0.59 0.64 0.62 0.62 0.54 0.66
so 0.71 0.65 0.64 0.69 0.72 0.63
rw 0.68 0.72 0.72 0.72 0.71 0.74
tw 0.67 0.64 0.66 0.65 0.64 0.67
ln 0.54 0.56 0.52 0.54 0.53 0.56
wo 0.50 0.41 0.42 0.51 0.45 0.41
om 0.54 0.50 0.54 0.58 0.57 0.56
ny 0.63 0.59 0.63 0.55 0.56 0.61
ee 0.57 0.57 0.62 0.59 0.58 0.59
bm 0.38 0.42 0.49 0.39 0.46 0.39
ts 0.49 0.54 0.46 0.52 0.52 0.50
fon 0.46 0.45 0.43 0.44 0.45 0.46
ti 0.38 0.28 0.34 0.36 0.37 0.34
sw 0.79 0.78 0.81 0.79 0.73 0.82
am 0.64 0.57 0.61 0.56 0.61 0.62
zu 0.63 0.66 0.68 0.58 0.64 0.68
ary 0.63 0.57 0.57 0.56 0.57 0.60
lg 0.47 0.56 0.52 0.52 0.51 0.51
tn 0.60 0.61 0.58 0.59 0.58 0.60
rn 0.45 0.48 0.46 0.52 0.48 0.51
ckb 0.80 0.82 0.81 0.81 0.81 0.81
ks 0.41 0.43 0.46 0.38 0.43 0.40
awa 0.53 0.46 0.53 0.52 0.51 0.52
ro 0.79 0.80 0.79 0.77 0.83 0.79
id 0.85 0.79 0.83 0.85 0.80 0.85
hu 0.81 0.83 0.82 0.85 0.81 0.85
nl 0.80 0.82 0.86 0.81 0.82 0.83
uk 0.82 0.83 0.81 0.83 0.84 0.83
yo 0.66 0.64 0.67 0.70 0.63 0.56
kk 0.81 0.82 0.82 0.82 0.83 0.80
mr 0.78 0.77 0.77 0.75 0.73 0.78
te 0.80 0.81 0.82 0.80 0.83 0.86
ht 0.68 0.66 0.71 0.69 0.65 0.70
nso 0.41 0.54 0.57 0.48 0.41 0.48
lij 0.55 0.55 0.46 0.57 0.54 0.52
ace 0.18 0.17 0.16 0.19 0.22 0.23
shn 0.71 0.74 0.74 0.69 0.73 0.74
mai 0.64 0.65 0.61 0.61 0.65 0.66
ban 0.59 0.66 0.64 0.68 0.64 0.66
mni 0.24 0.16 0.21 0.22 0.18 0.24
mi 0.56 0.54 0.55 0.51 0.53 0.52
pag 0.44 0.39 0.39 0.40 0.44 0.47
vi 0.82 0.84 0.82 0.83 0.83 0.84
war 0.46 0.54 0.56 0.47 0.47 0.59
tt 0.74 0.72 0.77 0.69 0.73 0.69
azb 0.49 0.53 0.48 0.42 0.50 0.47
ceb 0.46 0.50 0.49 0.47 0.52 0.53
min 0.54 0.51 0.47 0.48 0.49 0.49
Table 8: Fine-tuning results for monolingual models on SIB200. Languages are segmented by tier.
Masakhane
Lang absolute relative entropy
low high low high low high
MasakhaNER 2
pcm 0.75 0.74 0.74 0.75 0.74 0.74
ha 0.80 0.79 0.81 0.81 0.82 0.82
ig 0.85 0.84 0.84 0.85 0.84 0.84
sn 0.90 0.91 0.90 0.91 0.91 0.92
rw 0.77 0.76 0.78 0.78 0.76 0.76
tw 0.74 0.74 0.74 0.74 0.74 0.74
wo 0.74 0.69 0.72 0.74 0.73 0.71
ny 0.86 0.86 0.87 0.86 0.86 0.86
ee 0.79 0.78 0.79 0.79 0.79 0.79
bm 0.72 0.68 0.72 0.72 0.71 0.72
fon 0.73 0.71 0.71 0.71 0.74 0.74
sw 0.90 0.90 0.90 0.90 0.90 0.90
zu 0.82 0.80 0.82 0.80 0.80 0.82
lg 0.82 0.83 0.82 0.84 0.83 0.83
tn 0.77 0.78 0.77 0.78 0.76 0.78
yo 0.83 0.84 0.83 0.84 0.83 0.84
MasakhaNEWS
pcm 0.88 0.90 0.90 0.88 0.89 0.89
ha 0.86 0.88 0.87 0.85 0.86 0.87
ig 0.86 0.84 0.86 0.83 0.86 0.83
sn 0.90 0.89 0.88 0.89 0.87 0.90
so 0.69 0.68 0.70 0.69 0.68 0.69
ln 0.77 0.79 0.81 0.81 0.77 0.79
om 0.79 0.79 0.80 0.80 0.78 0.80
ti 0.63 0.59 0.61 0.63 0.62 0.60
sw 0.80 0.81 0.80 0.82 0.83 0.80
am 0.86 0.86 0.86 0.86 0.85 0.86
lg 0.86 0.86 0.88 0.85 0.88 0.86
rn 0.71 0.72 0.69 0.72 0.70 0.69
yo 0.84 0.85 0.85 0.87 0.86 0.86
AfriSenti-Twitter
pcm 0.60 0.60 0.59 0.60 0.60 0.60
ha 0.74 0.73 0.73 0.72 0.70 0.73
ig 0.74 0.75 0.74 0.75 0.74 0.73
rw 0.59 0.59 0.59 0.59 0.59 0.60
tw 0.64 0.62 0.63 0.65 0.63 0.63
ts 0.48 0.50 0.54 0.50 0.53 0.50
sw 0.54 0.55 0.58 0.56 0.59 0.57
am 0.40 0.43 0.48 0.32 0.28 0.14
ary 0.40 0.40 0.39 0.42 0.40 0.41
yo 0.67 0.68 0.67 0.67 0.67 0.69
Table 9: Fine-tuning results for monolingual models on MasakhaNER, MasakhaNEWS and AfriSenti-Twitter. Languages are segmented by tier.
Appendix C Thresholding Details

### v2

@@ Appendix A Training Hyperparameters and Finetuning Results
Table 12: XLM-R base, XLM-R with continued pretraining on unseen languages with raw_wiki, +primary and +heuristic partitions fine-tuned on SIB200. Results are grouped by tier.
XLMR Continued Pretraining (Belebele)
wiki baseline raw_wiki +primary +heuristic
wo 0.27 0.26 0.29 0.26
ts 0.24 0.25 0.33 0.25
ny 0.26 0.25 0.27 0.25
rw 0.24 0.23 0.30 0.27
ig 0.26 0.25 0.30 0.22
ln 0.25 0.24 0.30 0.24
sn 0.25 0.23 0.33 0.22
bm 0.24 0.26 0.24 0.25
ti 0.24 0.27 0.29 0.24
lg 0.24 0.24 0.28 0.26
tn 0.25 0.24 0.33 0.25
zu 0.25 0.23 0.32 0.27
shn 0.25 0.26 0.27 0.24
nso 0.25 0.26 0.31 0.26
ht 0.24 0.24 0.33 0.25
yo 0.26 0.25 0.27 0.23
mi 0.26 0.27 0.29 0.24
Avg 0.25 0.25 0.30 0.25
Table 13: XLM-R base, XLM-R with continued pretraining on unseen languages with raw_wiki, +primary and +heuristic partitions fine-tuned on MC and tested on belebele. Languages are grouped by tier.
Multilingual Pretraining (XNLI)
wiki raw_wiki +primary +heuristic
el 0.39 0.39 0.39
bg 0.42 0.40 0.38
de 0.46 0.43 0.43
tr 0.43 0.41 0.41
th 0.38 0.37 0.38
fr 0.48 0.46 0.47
ru 0.40 0.38 0.37
es 0.50 0.49 0.50
zh 0.43 0.43 0.39
en 0.72 0.71 0.70
ar 0.39 0.38 0.38
sw 0.41 0.39 0.40
ur 0.38 0.39 0.37
hi 0.38 0.37 0.36
vi 0.41 0.41 0.41
Avg 0.44 0.43 0.42
Table 14: Multilingual models pretrained on top 100 Wikipedias with raw_wiki, +primary and +heuristic fine-tuned on MNLI and tested on XNLI. Languages are grouped by tier.
Multilingual Pretraining (Belebele)
wiki raw_wiki +primary +heuristic
pa 0.27 0.26 0.27
ta 0.24 0.23 0.25
af 0.24 0.23 0.25
th 0.22 0.25 0.24
fr 0.24 0.26 0.25
es 0.22 0.26 0.24
cs 0.26 0.24 0.25
pl 0.26 0.24 0.25
de 0.27 0.26 0.26
en 0.25 0.26 0.25
fi 0.26 0.25 0.26
pt 0.25 0.27 0.25
hr 0.25 0.27 0.27
el 0.24 0.23 0.23
da 0.26 0.25 0.24
it 0.25 0.26 0.24
bn 0.25 0.27 0.25
ml 0.27 0.24 0.25
is 0.25 0.27 0.27
bg 0.27 0.24 0.25
kn 0.26 0.25 0.25
et 0.27 0.26 0.25
ca 0.28 0.26 0.28
ko 0.27 0.20 0.25
tr 0.24 0.26 0.23
he 0.27 0.24 0.27
zh 0.25 0.24 0.24
sl 0.25 0.25 0.26
ja 0.27 0.24 0.23
lt 0.26 0.23 0.24
sk 0.23 0.24 0.25
ky 0.24 0.26 0.26
ka 0.25 0.25 0.26
mk 0.27 0.25 0.26
uk 0.24 0.26 0.23
hu 0.25 0.27 0.27
ro 0.25 0.27 0.24
id 0.26 0.27 0.27
hi 0.26 0.24 0.27
jv 0.23 0.24 0.29
hy 0.26 0.24 0.23
my 0.23 0.24 0.23
nl 0.26 0.23 0.24
sv 0.22 0.26 0.25
yo 0.23 0.26 0.25
vi 0.24 0.26 0.22
tg 0.24 0.24 0.22
su 0.24 0.28 0.26
gu 0.23 0.24 0.24
eu 0.26 0.25 0.25
ht 0.26 0.26 0.25
kk 0.27 0.26 0.27
mr 0.28 0.24 0.23
ceb 0.25 0.26 0.26
war 0.24 0.23 0.26
Avg 0.25 0.25 0.25
Table 15: Multilingual models pretrained on top 100 Wikipedias with raw_wiki, +primary and +heuristic fine-tuned on MC and tested on belebele. Languages are grouped by tier.
Appendix B Thresholding Details